OpenStudy (anonymous):

Two triangles can be formed with the given information. Use the Law of Sines to solve the triangles. B = 29°, b = 26, c = 28

OpenStudy (anonymous):

two answers are possible .because the congruency is not sure a,s,sproperty will not give unique answer. but any way you use sine formula to solve

OpenStudy (anonymous):

A = 92.5°, C = 58.5°, a = 53.6; A = 87.5°, C = 121.5°, a = 53.6 A = 92.5°, C = 58.5°, a = 12.6; A = 87.5°, C = 121.5°, a = 12.6 A = 119.5°, C = 31.5°, a = 46.7; A = 2.5°, C = 148.5°, a = 2.3 A = 119.5°, C = 31.5°, a = 14.5; A = 2.5°, C = 148.5°, a = 14.5

OpenStudy (anonymous):

sinA/a=sinB/b=sinC/c

OpenStudy (anonymous):

im not clear on how to do that though:/

OpenStudy (anonymous):

sinB/b=sinC/c solve for C

OpenStudy (anonymous):

use cosine formula to solve for a.namely a^2=b^2+c^2-2bc cosA

OpenStudy (anonymous):

here A will have two values according to A+B+C=180. Amay be accute or obtuse

OpenStudy (anonymous):

ok so how would i plug in those numbers to the cos formula? im not good at this at all

OpenStudy (anonymous):

i just dont know where the A comes from

OpenStudy (anonymous):

first B may have two values as sin function is positive in 1st and 2nd quadrent. for both the values you solve for A.If B+C is more than 180 Bcan not be obtuse. sssthere is only one solution.

OpenStudy (anonymous):

triangle prperty .Sum of aaaaangles in a triangle is 180 deg

OpenStudy (anonymous):

ok so how can i find the answer? im still lost

OpenStudy (anonymous):

first step use sinB/b=sinC/c

OpenStudy (anonymous):

find sinC

OpenStudy (anonymous):

i did and i got a weird decimal. how do i find sinC

OpenStudy (anonymous):

sinB/b=.018

OpenStudy (anonymous):

namely sinC=(c/b)sinB

OpenStudy (anonymous):

WHAT IS YOUR SINc

OpenStudy (anonymous):

sinC=.522?

OpenStudy (anonymous):

REFER THE TRIGNOMETRY TABLE AND FIND THE ANGLE FOR WHICH .522 IS OCCURING

OpenStudy (anonymous):

i dont see it on there